6. General Electric Incorporated issued a 30 year zero-coupon bond. If comparable AA rated bonds yield 7.8%, what is the price of bond? (Discount at an annual rate)

(a) $1,000.00 (b) $       0.00 (c) $1,050.60 (d) $   105.06 (e) $   780.00

7. A bond with a bond rating of BBB or higher by Standard and Poor's, or Baa or higher by Moody's is referred to as being what type of bond

(a) investment grade (b) subordinated (c) debenture (d) mortgage (e) junk

8. The 10 year IBM corporate bond you purchased is contracted to pay a semi-annual 6.75% coupon. If you purchase this bond for $1,089.64 it will have a yield to maturity of 5.569%. What is the expected capital gain or loss for the first year holding period?

(a) a capital loss of 0.62% (b) a capital loss of 1.18% (c) a capital loss of 3.40% (d) a capital gain of 1.18% (e) a capital gain of 0.62%
Answer #1

6. Price of Bond =Par Value/(1+YTM)^30 =1000/(1+7.8%)^30 =105.06 (option d is correct option)

7. If bond rating is greater than BBB then it is option a Investent grade.

8.Current Yield =Coupon/Price =67.5/1089.64 =6.19%
Capital loss =YTM-Current Yield =5.569%-6.19% =-0.62% (Capital loss of 0.62%)